OpenStudy (anonymous):

okay, so i will give a medal to whoever can sucsessfully teach me how to solve this.... 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

i dint know

OpenStudy (anonymous):

@kevin_hart

OpenStudy (anonymous):

its solving system of equations using substittion

OpenStudy (anonymous):

@dmezzullo please help??

OpenStudy (anonymous):

Interesting... 2x + y = 1 4x + 2y = -2 multiplying "2x + y = 1" by 2 gives you: 4x+2y = 2 so you have: 4x + 2y = 2 4x + 2y = -2 So the same equation is +/-2 FWIW...

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

(solving for x) 2x+y=1 (subtract y) 2x-y=1-y 2x=-y+1 (no here you will divide by two) 2x/2= (-y+1)/2 so x=(-y+1)/2 once you have found x plug (-y+1)/2 on either equation (remember just for x) so 2x + y = 1 2((-y+1)/2) + y + 1 -------> Solve this ^^^^ will give you Y once you have Y just plug y into any of the original functions and so then you find x

OpenStudy (anonymous):

You mean that actually helped? I was sitting here wondering what I was supposed to do with it!!! :-) Anyway, glad to be of assistance! :-) And thank you!